OpenStudy (anonymous):

Find d2y/dx2 by implicit differentiation, given x^2 + 3y^2 = 5. Help please!

OpenStudy (anonymous):

step number 1 take the derivative wrt x you get \[2x+6yy'=0\] is that part ok?

OpenStudy (anonymous):

yes I got that

OpenStudy (anonymous):

then solve for \[y'\] via \[6yy'=-2x\] \[y'=-\frac{x}{3y}\]

OpenStudy (anonymous):

Mhm I got that far

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

ok so now we do it again, this time with the paininthebutt quotient rule

OpenStudy (anonymous):

OOOOH

OpenStudy (anonymous):

want to factor out the \[-\frac{1}{3}\] first or leave it in?

OpenStudy (anonymous):

we can leave it in I guess. Which ever one is easier...

OpenStudy (anonymous):

lets make life easy and factor it out so we have \[-\frac{1}{3}\times \frac{x}{y}\] and we can mutiply by \[-\frac{1}{3}\] at the end

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

so now for \[y''=-\frac{1}{3}\times \frac{y-xy'}{y^2}\]

OpenStudy (anonymous):

we are not done yet, but is that ok?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok and last step is an algebra one. because we actually know what \[y'\] is

OpenStudy (anonymous):

\[y'=-\frac{x}{3y}\]so that is what we replace it by in our second derivative to get \[y''=-\frac{1}{3}\times \frac{y-x\times -\frac{x}{3y}}{y^2}\]

OpenStudy (anonymous):

still not done. more algebra

OpenStudy (anonymous):

ok >.<

OpenStudy (anonymous):

need to clear the compound fraction, so maybe multply top and bottom by \[3y\] to get \[y''=-\frac{3y^2+x^2}{9y^3}\]

OpenStudy (anonymous):

and we are STILL NOT DONE

OpenStudy (anonymous):

because in a fit of trickery they have arranged it so that you numerator is in fact what you started with (be on the lookout for this) recall that we started with \[x^2+3y^2=5\] so the numerator is just... 5! and you get \[y''=-\frac{5}{9y^3}\] (check my algebra)

OpenStudy (anonymous):

You... are a genius. I was about to cry. Thank you sooo much!

OpenStudy (anonymous):

actually myininay will check my algebra, she loves finding my mistakes

OpenStudy (anonymous):

ooops i got a medal. i guess it is right

OpenStudy (anonymous):

You were completely correct! Thank you again!

myininaya (myininaya):

well there is one thing i don't like that you said...

OpenStudy (anonymous):

yw

myininaya (myininaya):

i just don't like \[\times \] this symbol

myininaya (myininaya):

but goof job

myininaya (myininaya):

i mean good job lol

OpenStudy (anonymous):

oh i thought you were going to tell me i was off by a minus sign or something. i put that in to make sure it is clear that we were multiplying. i like "goof job" better actually

myininaya (myininaya):

me too